OpenStudy (narissa):

simplify 7x +4y+3(x-y)-2x

OpenStudy (narissa):

i have to justify my answers

OpenStudy (fibonaccichick666):

so what have you done so far?

OpenStudy (narissa):

2x

OpenStudy (fibonaccichick666):

uh, that doesn't make sense, sorry

OpenStudy (narissa):

7x+3x-2x=8x

OpenStudy (fibonaccichick666):

so that is correct if only working with the x's, but probably not what your teach wants

OpenStudy (narissa):

no thats what i have so far

OpenStudy (fibonaccichick666):

typically for this, they want a step by step with which property you are using that allows you to do a specific action

OpenStudy (fibonaccichick666):

let me give you an example and it may help. 1 moment

OpenStudy (narissa):

+4y+3y=7y

OpenStudy (fibonaccichick666):

Problem: \(3x+2y-4(x+y)-7y\) Steps: 1. First I will use the distributive property to simplify -4(x+y). \[3x+2y-4(x+y)-7y=3x+2y-4x-4y-7y\] 2. Next I use the commutative property to group like terms \[3x+2y-4x-4y-7y=3x-4x+2y-4y-7y\] 3. Finally I combine the like terms and then use the distributive property to get rid of some negative signs \[3x-4x+2y-4y-7y=-1x-9y=-(x+9y)\]

OpenStudy (fibonaccichick666):

So the is the type of solution your teacher is most likely looking for. Can you attempt your problem in this way?

OpenStudy (narissa):

im trying give me a sec

OpenStudy (fibonaccichick666):

please do, feel free to upload a photo if you do not wish to type it all out

OpenStudy (narissa):

-1x+9y

OpenStudy (narissa):

k im stuck

OpenStudy (narissa):

help what do i do next?

OpenStudy (fibonaccichick666):

huh?

OpenStudy (fibonaccichick666):

sorry I don't know what you did. Not your problem

OpenStudy (fibonaccichick666):

that was my example

OpenStudy (narissa):

oh lol now im really confused

OpenStudy (fibonaccichick666):

oh oh wait a sec. Wow I made a similar ex

OpenStudy (fibonaccichick666):

no i didn't, so you are supposed to do your problem the same way I did my example above. Write out every step like I did

OpenStudy (narissa):

ok thanks

OpenStudy (fibonaccichick666):

only with your posed question

OpenStudy (fibonaccichick666):

you will not need the final distributive step I did though

OpenStudy (narissa):

im totally confused thanks anyway

OpenStudy (narissa):

@satellite73

OpenStudy (fibonaccichick666):

where, i'm sorry, But could you use full sentences to explain why you are confused so I can assist?

OpenStudy (narissa):

@jim_thompson5910

OpenStudy (narissa):

@dan815

OpenStudy (narissa):

@Nnesha

OpenStudy (narissa):

@Loser66

OpenStudy (fibonaccichick666):

Well here is my example again. Example problem: Problem: Simplify \(3x+2y−4(x+y)−7y\) Steps: 1. First I will use the distributive property to simplify \[-4(x+y). 3x+2y−4(x+y)−7y=3x+2y−4x−4y−7y\] 2. Next I use the commutative property to group like terms \[3x+2y−4x−4y−7y=3x−4x+2y−4y−7y\] 3. Finally I combine the like terms and then use the distributive property to get rid of some negative signs \[3x−4x+2y−4y−7y=−1x−9y=−(x+9y)\]

OpenStudy (narissa):

7x+3x-2x+4y-3y

OpenStudy (narissa):

8x+1y

OpenStudy (narissa):

is that how i do it?

OpenStudy (fibonaccichick666):

you skipped a bunch of steps and did not explain each

OpenStudy (fibonaccichick666):

you must show every single minute step AND give full explanation in order to get full credit I would assume

OpenStudy (narissa):

yes but its hard im in a special class for math and reading

OpenStudy (fibonaccichick666):

I fail to see why. You are clearly performing the steps, just not writing them down

OpenStudy (fibonaccichick666):

literally follow the example I wrote out. That is the format of answer they are looking for

OpenStudy (fibonaccichick666):

Does the example make sense?

OpenStudy (fibonaccichick666):

I mean you are getting the correct answer.

OpenStudy (narissa):

yes but its hard im going to try again

OpenStudy (fibonaccichick666):

just not showing your work as requested

OpenStudy (fibonaccichick666):

What is hard? (not meant to be rude, but you have done the problem correctly. I fail to understand which part is hard.)

OpenStudy (narissa):

to write all the steps

OpenStudy (fibonaccichick666):

(again, nit to be rude, but) why?

OpenStudy (fibonaccichick666):

I am honestly curious. (I teach and I wish to fully understand the difficulty)

OpenStudy (narissa):

k does the first step ;lool like this

OpenStudy (fibonaccichick666):

Maybe try a question approach. How did you go from \[7x +4y+3(x-y)-2x\rightarrow 7x+3x-2x+4y-3y\]

OpenStudy (narissa):

because its easier to think and solve the problem all in my head

OpenStudy (narissa):

ok should the first step look like this----3(x-y)*7x+4y+3(x-y)-2x=7x+4y+3x-3y-2x

OpenStudy (fibonaccichick666):

but you are doing it in your head, why can't you write it? (i realize it seems redundant, but it's not like you are just having an epiphany to go from step to step) Not to be offensive, do you have autism? I have only met one student with that issue and that was the reason for him.

OpenStudy (narissa):

yes and a.d.h.d

OpenStudy (fibonaccichick666):

no, because that left side will not give you that right side. And ah, thank you. Let me change my approach then

OpenStudy (fibonaccichick666):

ok, so are you able to read my template?

OpenStudy (fibonaccichick666):

(your right side up there is actually step one, but you just write the initial problem for the left)

OpenStudy (narissa):

please do that i need to get 2 more done if i upload the paper and tell u all the answers can you just tell me if they are right im failing in math,english and biology and report cards are coming out in 9 days

OpenStudy (fibonaccichick666):

I'm sorry, but chances are (in math) the reason your answers are wrong are due to lack of work :/

OpenStudy (narissa):

k thats just so confusing i will be back in 4 minutes got to take a quick break ive been on here almost all day

OpenStudy (fibonaccichick666):

So you probably have the correct answer, but it is the process they are grading.

OpenStudy (narissa):

because i do not know how to explain

OpenStudy (fibonaccichick666):

so can you copy my example and use it as a template?

OpenStudy (fibonaccichick666):

like literally copy it, then we can go through

OpenStudy (narissa):

i did i wrote it all down

OpenStudy (fibonaccichick666):

ok so, step one, try my template out, replace my problem with yours to start

OpenStudy (fibonaccichick666):

step two, try and see what happened in each step. I know it seems really really stupid to write everything out like that, but that is the 'process' the teacher wants

OpenStudy (fibonaccichick666):

From what I have learned about autism, learning a template will probably help. So if you memorize the example and then we make it generic, you will be able to do any similar problem. essentially the teacher just wants you to answer questions. What did I do? Why did I do it?

OpenStudy (narissa):

be back in 2 minutes dont leave please

OpenStudy (fibonaccichick666):

also side note... I hate common core. This unit is so not mathematically logical

OpenStudy (fibonaccichick666):

ok, I'll be back in 5 break time :)

OpenStudy (narissa):

thanks

OpenStudy (fibonaccichick666):

alrighty

OpenStudy (fibonaccichick666):

so, I have an analogy for you.

OpenStudy (narissa):

k im back still trying to do this

OpenStudy (fibonaccichick666):

have you ever put a puzzle together?

OpenStudy (narissa):

no i cant unless its small

OpenStudy (fibonaccichick666):

small is fine, let me find a better analogy. What do you like? What are you good at?

OpenStudy (fibonaccichick666):

do you like trains?

OpenStudy (narissa):

playing video games

OpenStudy (fibonaccichick666):

ok, which games?

OpenStudy (narissa):

call of duty black ops 2

OpenStudy (narissa):

k look i think i got the first step trying to do the second

OpenStudy (fibonaccichick666):

ohk, so in black ops, you have to play the story line piece by piece right?

OpenStudy (narissa):

7x+4y+3(x-y)=7x+4y+3x+3y-2x

OpenStudy (fibonaccichick666):

you can't move on till you beat one mission(if you are like me, until you beat it on veteran :P )

OpenStudy (fibonaccichick666):

very very close, you forgot the -2x on the left

OpenStudy (narissa):

not really its a shooting game with friends online oh well they call them my friends

OpenStudy (narissa):

oops yaaa! i wrote that but didn't type it now im working on the nxt

OpenStudy (fibonaccichick666):

good, now before you move on

OpenStudy (fibonaccichick666):

what property did you use to simplify 3(x-y)?

OpenStudy (fibonaccichick666):

also, you dropped the minus sign(but I know it is a type-o since you did it right above)

OpenStudy (narissa):

distributive?

OpenStudy (fibonaccichick666):

yep. and that my dear, is your justification